Is the rate of hay fever lower for people over the age of 50 as compared to people under the age of 25? A random sample of

0 0
Answer #1

Ans:

a)level of significance=0.05

H_0:\mu_1\leq \mu _2

H_a:\mu_1>\mu _2

b)We will use t-statistic.

pooled standard deviation=SQRT(((16-1)*15.41^2+(14-1)*11.57^2)/(16+14-2))=13.761

standard error=13.761*sqrt((1/16)+(1/14))=5.036

Test statistic:

t=(109.5-99.36)/5.036

t=2.014

c)df=16+14-2=28

p-value=tdist(2.014,28,1)=0.0269

d)As,p-value<0.05,we reject the null hypothesis.

Yes,data is statistically significant at level